OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (25, 30) and (125, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points?

OpenStudy (anonymous):

Find slope I guess: y2 - y1 / x2 - x1 50 - 30 = 20 / 125 - 25 = 100 Slope: 20/5 simplified: 4/1 = 4 So it would have a slope of 4 or 4/1

OpenStudy (anonymous):

so you had to find the x ?

OpenStudy (anonymous):

Well the question asks: what is the slope of the line between the two points? There is a lot of extra information in this question that could get you confused but all you need to focus on is the main question:)

OpenStudy (anonymous):

So I wasn't finding x i was finding the slope of the line